OpenStudy (anonymous):

Find the value of the variable. If your answer is not an integer, leave it in simplest radical form. Answer A. 5√ 2 B. 5√ 2/3 C. 5√ 2/2 D. 5√ 3

OpenStudy (anonymous):



OpenStudy (zzr0ck3r):

sin(45)/x = sin(90)/5

OpenStudy (anonymous):

sin of 45 is .7071

OpenStudy (zzr0ck3r):

5sin(45)/sin(90) = x 5(sqrt(2)/2)/sin(90) = x 5sqrt(2)/2 = x

OpenStudy (anonymous):

thanks

OpenStudy (zzr0ck3r):

dont aprx untill the end

OpenStudy (anonymous):

so the answer is C

OpenStudy (zzr0ck3r):

o I thought it said aprx...dont aprx at all:) yes c

OpenStudy (anonymous):

ok thanks

OpenStudy (zzr0ck3r):

np is this how you have been solving them? law of sines